OpenStudy (anonymous):

find the slope. 4y=12x+36. PLEASE HELP

OpenStudy (dmndlife24):

divide by 4 on both sides y = 3x + 9 And since m = slope in y = mx +b 3 = slope

OpenStudy (anonymous):

Hello, the general form of the function is \(y=ax+b\), so \(a\) will be the slope.

OpenStudy (anonymous):

YOU HAVE TO SOLVE THE EQUATION. @tinybookworm

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

SO HOW IS THE SLOPE GOING TO BE ''a''

OpenStudy (anonymous):

THANKS @dmndlife24

OpenStudy (anonymous):

\(a\) is the number before x. For example, in the function \(y=2x\), the slope is 2. Or in the function \(y=-5x-7\), the slop will be -5. Like that

OpenStudy (anonymous):

OH K. THANKS

OpenStudy (anonymous):

You are welcome

OpenStudy (dmndlife24):

Happy to help